Circular No. 8986

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 S2 = C/2001 D1 = C/2004 X7 (SOHO) Further to IAUC 8985, an additional near-sun presumed comet has

been found by R. Kracht on SOHO website images (discovery observation tabulated below). C/2008 S2 has been identified by Kracht as apparently identical with C/2001 D1 (cf. IAUC 7936) and C/2004 X7 (cf. IAUC 8466), a linkage that has been verified by B. G. Marsden (see MPEC 2008-S81 and 2008-S82). K. Battams writes that C/2008 S2 was tiny and stellar in appearance (mag 7.5–8.0) and appeared even fainter on the far side of the sun. In remeasuring the positions of C/2001 D1, Battams notes that it was then very faint and tiny (mag ∼ 8), and was slightly brighter when visible on the far side of the sun.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 S2 Sept. 17.814 11^{h}50.^{m}0 + 1^{o}36′ C2 RK 2008-S81

COMET C/2008 S3 (BOATTINI) A. Boattini reports his discovery of a comet on CCD images obtained

with the 1.5-m reflector at Mt. Lemmon (discovery observation tabulated below), the comet having strong condensation with a coma ∼ 10′′ wide but elongated ∼ 15′′ at p.a. ≈ 115^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, D. T. Durig (Sewanee, TN, U.S.A., 0.30-m f/2.5 Schmidt-Cassegrain reflector, Sept. 30.4 UT) reports a tail ∼ 10′′ long in p.a. 90^{o}. E. Guido (Castellammare di Stabia, Italy) writes that G. Sostero, P. Camilleri, and he co-added fifty 60-s unfiltered exposures that were obtained remotely with a 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A., on Sept. 30.4, showing a coma of diameter ∼ 8′′ with a central condensation and slight elongated toward the northwest. J. Young (Table Mountain 0.61-m f/16 Cassegrain reflector, Sept. 30.5) finds a bright, diffuse coma of diameter 6′′ with a very short tail ∼ 16′′ long in p.a. 140^{o}–160^{o}, curving to the east.

2008 UT  α_{2000}  δ_{2000}  Mag.
Sept.29.49105  4^{h}42^{m}49.^{s}71  +18^{o}29′48′′.3  18.4

The available astrometry, the very uncertain parabolic orbital elements (T = 2011 Apr. 4.172 TT, q = 3.12838 AU, ω = 95^{o}.199, Ω = 56^{o}.737, i = 157^{o}.926, equinox 2000.0), and an ephemeris appear on MPEC 2008-S96.

2008 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green